FACTORIZATION OF PIECEWISE CONSTANT MATRIX FUNCTIONS AND SYSTEMS OF LINEAR DIFFERENTIAL

EQUATIONS

TORSTEN EHRHARDT ILYA M. SPITKOVSKY

Abstract. Let G be a piecewise constant n × n matrix function which is defined on a smooth closed curve Γ in the complex sphere and which has m jumps. We consider the problem of determining the partial indices of the factorization of the matrix function G in the space L^{p}(Γ). We show that this problem can be reduced to a certain problem for systems of linear differential equations.

Studying this related problem, we obtain some results for the partial indices for general n and m. A complete answer is given for n = 2, m = 4 and for n = m = 3. One has to distinguish several cases. In some of these cases, the partial indices can be determined explicitly. In the remaining cases, one is led to two possibilities for the partial indices. The problem of deciding which is the correct possibility is equivalent to the description of the monodromy of n-th order linear Fuchsian differential equations with m singular points.